OpenStudy (anonymous):

HELP ME!!!THANK YOU! GOD BLESS YOU! Find the inverse of the function f(x)=1/3 (x-4) f^-1(x)=_____

OpenStudy (anonymous):

Help!

OpenStudy (anonymous):

@Destinymasha @poopsiedoodle @nickmifsud1223 @marylou004 @camerondoherty @caitlinnr14 @CloverRacer @dan815 @DrPepperx3 @FibonacciChick666

OpenStudy (anonymous):

@kropot72 @Whitemonsterbunny17 @adrynicoleb @bibby

OpenStudy (anonymous):

@arabpride @timaashorty @triciaal @kirbykirby @inkyvoyd @perl

OpenStudy (anonymous):

Is it \[\large f(x) = \frac1{3(x-4)}\] or\[\large f(x) =\frac13(x-4)\] ?

OpenStudy (anonymous):

The second one

OpenStudy (anonymous):

Okay. Here's how to do it ^^ First, replace f(x) with y. \[\large y = \frac13(x-4)\] NEXT, switch x and y. \[\large x = \frac13(y-4)\] Now just solve for y :)

OpenStudy (anonymous):

How do I solve for y?

OpenStudy (anonymous):

Well, if I might have an example... Say you have \[\large x = \frac1{3y + 8}\] First, multiply 3y+8 to both sides... \[\Large x(3y+8) = \frac1{3y+8}\cdot (3y+8)\] \[\Large x(3y+8) = \frac1{\cancel{3y+8}}\cdot \cancel{(3y+8)}\] \[\Large x (3y+8) = 1\] Now, divide both sides by x. \[\Large 3y+8 = \frac1x\] Next, subtract 8 from both sides to get \[\Large 3y = \frac1x-8\] \[\Large 3y = \frac{1-8x}{x}\] Finally, divide both sides by 3. \[\Large y = \frac{1-8x}{3x}\] Mostly algebraic manipulation :D